OpenStudy (anonymous):

Solve: 4(x-1)= 2(x+2)+6

OpenStudy (anonymous):

4x-4=2x+10 +4 +4 4x=2x+14 /2x /2x 2/2x=14/2 x=7

OpenStudy (anonymous):

thank you!(: